OpenStudy (anonymous):

(9m)^-2 Simplify. Using positive exponents only. Help please?

OpenStudy (lgbasallote):

9m - 2 is already in simplest form....

OpenStudy (anonymous):

well it's actualy (9m)^2

OpenStudy (anonymous):

sorry i mean (9m)^-2

OpenStudy (anonymous):

its asking me to use postive exponents only. I am confused.

OpenStudy (lgbasallote):

here's a hint \[\huge a^{-m} \implies \frac 1{a^m}\] does that help?

OpenStudy (anonymous):

is it 18m?

OpenStudy (anonymous):

i mean 81m

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (anonymous):

exponents really confuse me alot

OpenStudy (lgbasallote):

let me give an example \[\huge 2^{-2} \implies \frac 1{2^2} \implies \frac 14\]

OpenStudy (lgbasallote):

do you get it now?

OpenStudy (anonymous):

9m*2? am I on the right track?

OpenStudy (lgbasallote):

no...

OpenStudy (anonymous):

:(

OpenStudy (lgbasallote):

let me give another example \[\huge (3x)^{-2} \implies \frac 1{(3x)^2} \implies \frac 1{9x^2}\]

OpenStudy (anonymous):

would you mind showing me the steps? I am a first year algebra student and have no idea what I am supposed to be doing. Sorry that I am not getting your examples.

OpenStudy (lgbasallote):

like i said before...if you have a negative exponent..just put it in the denominator for example (3x)^(-2) to make the exponent positive, all you have to do is put it in the denominator like so \[\frac{1}{(3x)^2}\] follow?

OpenStudy (anonymous):

sort of, yes.

OpenStudy (anonymous):

so 1/(9m)^2 = 1/9m^2?

OpenStudy (lgbasallote):

no.. you have to square 9 as well

OpenStudy (lgbasallote):

see what i did earlier \[\frac 1{(3x)^2} \implies \frac 1{3^2 x^2} \implies \frac 1{9x^2}\]